# MuxServe: Flexible Multiplexing for Efficient Multiple LLM Serving

### v1

@@ Abstract
Large language models (LLMs) have demonstrated remarkable performance, and organizations are racing to serve LLMs of varying sizes as endpoints for use-cases like chat, programming and search. However, efficiently serving multiple LLMs poses significant challenges for existing approaches due to varying popularity of LLMs. In the paper, we present MuxServe, a flexible spatial-temporal multiplexing system for efficient multiple LLM serving. The key insight behind is to colocate LLMs considering their popularity to multiplex memory resources, and leverage the characteristics of prefill and decoding phases to separate and flexibly colocate them to multiplex computation resources. MuxServe formally formulates the multiplexing problem, and proposes a novel placement algorithm and adaptive batch scheduling strategy to identify optimal colocations and maximize utilization. MuxServe designs a unified resource manager to enable flexible and efficient multiplexing. Evaluation results show that MuxServe can achieves up to 1.8×1.8$\times$1.8 × higher throughput or processes 2.9×2.9$\times$2.9 × more requests within $99\%$ SLO attainment.
Machine Learning, ICML
@@ 3.1 Problem Formulation
        $best\_mesh.u$ = make_unit($m$, $d$)
@@ 3.3 Maximize Intra-unit Throughput
where $t^{m}_{p},t^{m}_{d}$ and $l^{m}_{o}$ represent the prefill latency, decoding latency and average generation length of a batch requests with size $b^{m}$ for LLM $m$, respectively. This formulation is based on the observation that prefill phases of different LLMs are generally executed sequentially and decoding phases can be executed concurrently. Therefore, the latency of a batch requests is equal to the sum of all prefill phases of different LLMs and the decoding phases of the LLM. The prefill and decoding latency of different batches and request length can be profiled in advance. The average generation length can be estimated from requests history or specific dataset, ShareGPT (ShareGPT-Team, 2023) for instance. Given the request arrival rates, we can use binary search to find the batch size $b$ that can satisfy the traffic.
@@ 3.4 Resource Manager
After finding the optimal LLM units and determining the batching and scheduling strategy, MuxServe requires a new mechanism to support flexible and efficient spatial-temporal multiplexing of LLMs due to the following challenges: different prefill and decoding jobs need to flexibly share the computation resources, and share the weights and KV cache to reduce memory waste and fragmentation. To address these challenges, MuxServe proposes a unified resource manager for flexible and efficient multiplexing. Figure 4 shows the overview of GPU resource management in an LLM unit.
The parallel runtime manages computation resources of an LLM unit in the granularity of SM. MuxServe schedules prefill and decoding jobs from colocated LLMs with ADBS algorithm, then the parallel runtime dynamically assigns SMs to each job at runtime rather than statically allocating. The implementation is based on CUDA MPS (NVIDIA, 2022b). As illustrated in Figure 4, the SMs are all allocated to one job at step 1, and allocated to two jobs at step 2.
The prominent challenge lies in sharing the memory resources among different jobs is to reduce memory waste and fragmentation. LLM weight and KV cache consume a huge amount of memory and need to be shared among jobs. Furthermore, KV cache increases dynamically, and different LLMs posses varying sizes of KV cache due to differences in the number of attention heads, hidden layers, and hidden sizes.
To efficiently share memory resources, MuxServe divides them into three partitions. The first partition is a unified KV cache space enabled by our head-wise cache. Leveraging the observation that the size of each attention head is often consistent or has limited variations across different LLMs, for example LLaMAs (Touvron et al., 2023) and GPT-3 (Brown et al., 2020) all use 128. MuxServe allocates KV cache in head-wise granularity and accommodates the KV cache of different LLMs in a unified space to share the memory. To reduce redundancy, the second partition stores a single replica of LLM weights that can be shared among prefill and decoding jobs. The final partition reserves space for activation, which is utilized during inference runtime.
@@
4 Evaluation
MuxServe is built atop vLLM (Kwon et al., 2023), an efficient single LLM serving system based on PyTorch (Paszke et al., 2019), and utilizes NVIDIA MPS (NVIDIA, 2022b) to partition SM resources. In this section, we evaluate MuxServe on both synthetic and real workloads. We also perform ablation studies to verify the effectiveness of individual components.
4.1 Experimental Setup
Cluster. We conduct experiments on a 4 node cluster, each is equipped with 8 NVIDIA A100 (80GB) GPUs.
@@ 4.1 Experimental Setup
Metrics. We use the aggregated throughput as our evaluation metric since our target is to evaluate the GPU utilization. Since different LLMs have different arrival rates, we use the rate to compute a weighted average of throughput.
Baselines. We compare MuxServe with two baselines. The first is widely used spatial partitioning, which serves an LLM separately on a group of GPUs. We serve each LLM with vLLM (Kwon et al., 2023), a state-of-the-art serving framework. Another baseline is temporal multiplexing similar to AlpaServe (Li et al., 2023). Since AlpaServe does not support multiplexing of multiple LLMs, we implement this baseline by ourselves with our unified KV cache. For temporal multiplexing, we colocate LLMs with the placement optimized by our placement algorithm, schedule LLMs with FCFS in a temporal manner, and batch each LLM with continuous batching.
Figure 5: Throughput and SLO attainment on synthetic workloads.
@@ 4.2 End-to-End Results for Synthetic Workloads
Workloads. For synthetic workloads, we first generate request rates for each LLM using power-law distribution with an exponent $\alpha$, then generate requests arrival time with poisson processes. The requests are sampled from ShareGPT. We vary $\alpha$ and rate scales to evaluate a diverse workloads. For each $\alpha$, we first set the maximal request rate for each LLM to $20$ $req/s$, and then scale up the max rate and average rate for evaluation. The $\alpha$ decides the popularity of LLMs, and larger $\alpha$ means the fewer LLMs are more popular and receive a higher rates. Figure 8 shows the LLM traffic distribution as we vary $\alpha$. Typically, $\alpha=0.9$ or $2.1$ represent $20\%$ LLMs receives $50\%$ or $90\%$ request rates.
Results. Figure 5 shows the throughput and SLO attainment with varying $\alpha$ and average rates. The throughput of MuxServe outperforms two baselines in all scenarios, achieving up to 1.8×1.8$\times$1.8 × improvement. MuxServe can process up to 2.9×2.9$\times$2.9 × more requests within $99\%$ SLO attainment. When $\alpha$ is small, the request rates are more even and MuxServe can efficiently colocate prefill-decoding and decoding-decoding jobs to improve utilization. But the interference also brings some overhead, leading to a slightly lower SLO attainment with small SLO scale. With a larger $\alpha$, popular LLMs can be colocated with unpopular LLMs to multiplex memory resources, thus achieving a higher throughput with more SMs and larger KV caches. Popular LLMs can process more requests to achieve a higher SLO attainment.
@@ 6 Conclusion
7 Impact Statement
Large language models (LLMs) have been widely used in various areas, including code copilot, writing assistant, chatbots and enhancing search engine results. The broader impacts of LLMs have been extensively discussed, covering a wide range of considerations. While the various implications are recognized, no specific highlights need to be emphasized in this context.
Nevertheless, it is worth noting that serving LLMs can be both costly and energy-intensive, which in turn imposes limitations on their widespread adoption. Recognizing this challenge, the primary objective of this project is to develop more efficient methods for serving LLMs. By improving the efficiency of LLM serving, we aim to alleviate the associated costs and energy consumption, thereby facilitating broader and more sustainable usage of LLMs. This project can have several positive implications.
Increased accessibility: By reducing the resource requirements and energy consumption associated with serving LLMs, our methods can make LLM-based applications more accessible to a wider range of users, including those with limited computational resources or operating in energy-constrained environments.
Cost saving: Efficient LLM serving can lead to substantial cost savings, particularly for organizations or individuals relying heavily on LLM-based services. Lower resource demands translate into reduced infrastructure costs and operational expenses, making such services more economically viable.
Environmental sustainability: Given the energy-intensive nature of LLM serving, our research contributes to the reduction of carbon footprints and environmental impact. By optimizing resource utilization and minimizing energy consumption, we promote more sustainable practices in the deployment and operation of LLMs.
Enhanced productivity and user experience: By making LLM serving more efficient, our research has the potential to enhance productivity and user experience in various domains. Faster response times, reduced latency, and improved performance can lead to more effective code completion, better writing assistance, and more interactive and responsive chatbot interactions.
Overall, the broader impacts of our paper lie in advancing the state-of-the-art in LLM serving, making it more efficient, accessible, cost-effective, and environmentally sustainable. These advancements have the potential to positively influence various industries and domains relying on LLM-based technologies, ultimately benefiting end-users and promoting the wider adoption of LLM applications.

### v2

# MuxServe: Flexible Spatial-Temporal Multiplexing for Multiple LLM Serving

@@ Abstract
Large language models (LLMs) have demonstrated remarkable performance, and organizations are racing to serve LLMs of varying sizes as endpoints for use-cases like chat, programming and search. However, efficiently serving multiple LLMs poses significant challenges for existing approaches due to varying popularity of LLMs. In the paper, we present MuxServe, a flexible spatial-temporal multiplexing system for efficient multiple LLM serving. The key insight behind is to colocate LLMs considering their popularity to multiplex memory resources, and leverage the characteristics of prefill and decoding phases to separate and flexibly colocate them to multiplex computation resources. MuxServe formally formulates the multiplexing problem, and proposes a novel placement algorithm and adaptive batch scheduling strategy to identify optimal colocations and maximize utilization. MuxServe designs a unified resource manager to enable flexible and efficient multiplexing. Evaluation results show that MuxServe can achieves up to 1.8×1.8$\times$1.8 × higher throughput or processes 2.9×2.9$\times$2.9 × more requests within $99\%$ SLO attainment. The code is available at: https://github.com/hao-ai-lab/MuxServe.
Serving, Spatial-Temporal Multiplexing, Large Language Model
@@ 3.1 Problem Formulation
        $best\_mesh.u$ = make_unit($m$, $best\_mesh$)
@@ 3.3 Maximize Intra-unit Throughput
where $t^{m}_{p},t^{m}_{d}$ and $l^{m}_{o}$ represent the prefill latency, decoding latency and average generation length of a batch requests with size $b^{m}$ for LLM $m$, respectively. This formulation is based on the observation that prefill phases of different LLMs are generally executed sequentially and decoding phases can be executed concurrently, and different phases are interleaved. Therefore, the latency of a batch requests is equal to the sum of all prefill phases of different LLMs and the decoding phases of the LLM. The prefill and decoding latency of different batches and request length can be profiled in advance. The average generation length can be estimated from requests history or specific dataset, ShareGPT (ShareGPT-Team, 2023) for instance. Given the request arrival rates, we can use binary search to find the batch size $b$ that can satisfy the traffic. More details about the formulation please refers to Section A.2.
@@ 3.4 Resource Manager
After finding the optimal LLM units and determining the batching and scheduling strategy, MuxServe requires a new mechanism to support flexible and efficient spatial-temporal multiplexing of LLMs due to the following challenges: different prefill and decoding jobs need to flexibly share the computation resources, and share the weights and KV cache to reduce memory waste and fragmentation. To address these challenges, MuxServe proposes a unified resource manager for flexible and efficient multiplexing. Each LLM unit hosts a unified resource manager, and Figure 4 shows the overview of GPU resource management in an LLM unit.
The parallel runtime manages computation resources of an LLM unit in the granularity of SM based on NVIDIA MPS (NVIDIA, 2022b). MuxServe schedules prefill and decoding jobs from colocated LLMs with ADBS algorithm, then the parallel runtime dynamically assigns SMs to each job at runtime rather than statically allocating. This enables MuxServe to flexibly multiplex the execution of different LLMs. As illustrated in the right part of Figure 4, the SMs are all allocated to one job at step 1 due to its large computation intensity. After the execution, MuxServe schedules two jobs that can be executed concurrently at step 2 to share the SM resources.
The prominent challenge lies in sharing the memory resources among different jobs is to reduce memory waste and fragmentation. LLM weights and KV cache consume a huge amount of memory and need to be shared among jobs. Furthermore, KV cache increases dynamically, and different LLMs posses varying sizes of KV cache due to differences in the number of attention heads, hidden layers, and hidden sizes.
To efficiently share memory resources, MuxServe divides them into three partitions. The first partition is a unified KV cache space enabled by our head-wise cache. Leveraging the observation that the size of each attention head is often consistent or has limited variations across different LLMs, for example LLaMAs (Touvron et al., 2023) and GPT-3 (Brown et al., 2020) all use 128. MuxServe divides the KV cache table into small blocks, and each block holds the KV cache of one head for several tokens. This head-wise granularity enables MuxServe to accommodate the KV cache of different LLMs in a unified space to share the memory. To reduce redundancy, the second partition stores a single replica of LLM weights that can be shared among prefill and decoding jobs. The final partition reserves space for activation, which is utilized during inference runtime.
@@
Figure 5: Throughput and SLO attainment on synthetic workloads.
Implementation. MuxServe is built atop vLLM (Kwon et al., 2023), an efficient single LLM serving system based on PyTorch (Paszke et al., 2019), and utilizes NVIDIA MPS (NVIDIA, 2022b) to partition SM resources. The major component of MuxServe is implemented in a global scheduler, which runs on each LLM unit. The global scheduler manages a request queue and runs the ADBS algorithm to schedule requests from the queue. MuxServe disaggregates the prefill and decoding phases and launches separate vLLM processes, configured with different number of SM resources with NVIDIA MPS, as runtime engines for them. The global scheduler schedules prefill or decoding jobs to runtime engine processes via python mutiprocessing shared memory. MuxServe additionally runs a memory manager process to manage the unfied memory space. The runtime engines access to the memory space via CUDA IPC handler. When the runtime engine needs to allocate some KV cache for execution, it requests new cache blocks from the memory manager and fills the KV cache to the assigned blocks. After the execution of prefill jobs, the memory manager maintains the allocated KV cache blocks for later usage of corresponding decoding jobs. The KV cache will be released only when the request is finished. We implement the memory manager with C++ to optimize the overhead of block allocation.
@@
4 Evaluation
In this section, we evaluate MuxServe on both synthetic and real workloads. We also perform ablation studies to verify the effectiveness of individual components.
4.1 Experimental Setup
Cluster. We conduct experiments on a 4 node cluster, each is equipped with 8 NVIDIA A100 (80GB) GPUs. The intra-node connection is 600GB/s NVLink, and the inter-node connection is 200Gbps IB.
Metrics. We use the aggregated throughput as our evaluation metric since our target is to evaluate the GPU utilization. Since different LLMs have different arrival rates, we use the rate to compute a weighted average of throughput. We also evaluate the SLO attainment to study MuxServe’s effect on latency. SLO attainment measures the percentage of requests that can be finished within the latency target. In this paper, we scale the latency to different multiplies of single device execution latency (i.e. SLO scale).
Baselines. We compare MuxServe with two baselines. The first is widely used spatial partitioning, which serves each LLM separately on a group of GPUs. We serve each LLM with vLLM (Kwon et al., 2023), a state-of-the-art serving framework. Another baseline is temporal multiplexing similar to AlpaServe (Li et al., 2023). Since AlpaServe does not support multiplexing of multiple LLMs, we implement this baseline by ourselves with our unified KV cache. For temporal multiplexing, we colocate LLMs with the placement optimized by our placement algorithm, schedule LLMs with first-come-first-serve in a temporal manner, and batch the requests of each LLM with continuous batching.
@@ 4.2 End-to-End Results for Synthetic Workloads
Workloads. For synthetic workloads, we first generate request rates for each LLM using power-law distribution with an exponent $\alpha$, then sample the arrival time of each request with poisson processes. The requests are sampled from ShareGPT. We vary $\alpha$ and rate scales to evaluate a diverse workloads. For each $\alpha$, we first set the maximal request rate for each LLM to $20$ $req/s$, and then scale up the max rate and average rate for evaluation. The $\alpha$ decides the popularity of LLMs, and a larger $\alpha$ means that the fewer LLMs are more popular and receive a higher rates. Figure 8 shows the LLM traffic distribution as we vary $\alpha$. Typically, $\alpha=0.9$ or $2.1$ represent $20\%$ LLMs receives $50\%$ or $90\%$ of the total request rates.
Results. Figure 5 shows the throughput and SLO attainment with varying $\alpha$ and average rates. The throughput of MuxServe outperforms two baselines in all scenarios, achieving up to 1.8×1.8$\times$1.8 × improvement. MuxServe can also process up to 2.9×2.9$\times$2.9 × more requests within $99\%$ SLO attainment. When $\alpha$ is small, the request rates are more even and MuxServe can efficiently colocate prefill-decoding and decoding-decoding jobs to improve utilization. But the interference of colocation also brings some overhead, leading to a slightly lower SLO attainment with small SLO scale. With a larger $\alpha$, popular LLMs can be colocated with unpopular LLMs to multiplex memory resources, thus achieving a higher throughput with more SMs and larger KV caches. Popular LLMs can process more requests thus achieving a higher SLO attainment. The results also demonstrate that MuxServe can achieve much more significant improvement when the popularity among LLMs is more different (i.e. $\alpha$ is larger). The detailed results of P99 latency please refer to Appendix A.1.
@@ 4.4 Ablation Studies
The optimized placement also verifies our insight to prioritize the placement selection for LLMs with large computation requirements. In the 8 GPUs and 4 LLMs case, MuxServe colocates two popular small LLMs with one less popular small LLM on 4 GPUs, while allocating 4 GPUs for the less popular large LLM. In the 16 GPUs and 7 LLMs case, MuxServe colocates two large LLMs with low arrival rates on 4 GPUs, and splits the rest (majority) of GPUs into groups of 2 or 4 GPUs. MuxServe places the other 5 LLMs with high arrival rates on different groups considering their computation demands. To understand this, consider the contrary: if we prioritize the placement selection for LLMs with large memory requirements (i.e. try to balance memory consumption across GPUs). Large LLMs would be prioritized to be placed on large GPU groups without considering their popularity. Then small LLMs will be placed in the manner that tries to fit in the remaining GPU memory space. This is counterintuitive since popular small LLMs could need massive GPUs to fulfill the request traffic requirements.
@@ 6 Conclusion
Acknowledgements
We thank Chang Chen, Qianchao Zhu and Jinming Ma for their insightful review and discussions. The research is supported in part by Shanghai AI Laboratory, CUHK Interdisciplinary AI Research Institute, and the Centre for Perceptual and Interactive Intelligence (CPII) Ltd under the Innovation and Technology Commission (ITC)’s InnoHK.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Appendix
A.1 P99 Latency Comparison
Figure 12 shows the P99 average latency, TPOT (Time-Per-Output-Token) and TTFT (Time-To-First-Token) latency of three multiplexing approaches on synthetic workloads. The results demonstrate that MuxServe can achieve a lower P99 average latency compared with prior approaches. The P99 TPOT latency of MuxServe is a little bit higher than spatial partitioning due to interference, and significantly lower compared wth temporal multiplexing. The P99 TTFT latency of MuxServe is lower than both spatial partitioning and temporal multiplexing since MuxServe can significantly reduces queuing time. When $\alpha$ is large, the P99 TTFT latency is dominated by the most popular model. Temporal multiplexing can reduces the queuing time, thus obtaining a lower P99 TTFT compared with spatial partitioning.
Figure 11: The P99 average latency, TPOT (Time-Per-Output-Token) and TTFT (Time-To-First-Token) latency on synthetic workloads.
A.2 Throughput Estimator
Figure 12 shows an approximated execution timeline of MuxServe, which motivates the formulation of Equation 3. In a stable serving setting, requests arrive at a fixed interval. The prefill requests are executed one by one, with their decoding steps overlapping. Since the system is stable, it completes some requests every interval and receives new requests. Therefore, we can estimate the execution time of a stable batch $b^{m}$ as $\sum_{i\in b}t^{i}{p}+t^{m}{d}$. Hence we can estimate the throughput with Equation 3.
Figure 12: Execution timeline of MuxServe in a stable serving setting. $R.$ represents other requests in the batch.